Please answer in the box below
Use the following financial statements for Lake of Egypt Marina, Inc. LAKE OF EGYPT MARINA, INC. Balance Sheet as of December 31, 2018 and 2017 (in millions of dollars) 2018 2017 2018 2017 AssetS Current assets: Liabilities and Equity Current liabilities Cash and marketable securities 60 Accounts receivable s 40 50 s 32 40 28 50 140132 250$196 Accrued wages and taxes Accounts payable Notes payable Inventory 32 $140$104 s 11$124 50 Total Total Fixed assets tong-term debt: Gross plant and equipment Lesst Depreciation $300$220Stockholders equitys 7540Preferred stock (8 million shares)$8S8 Common atock and paid-in surplus Net plant and equipment Other long-term assets $225$180 (60 mil1ion shares) 60 60 24 Retained earnings Total Total assets $250 $204 $500--$400.. 104 $349$172 $400 Total Total liabilities and equity $500 LAKE OF EGYPT MARTNA, INC Income Statenent for Years Ending December 31, 2018 and 2017 (in millions of dollars) 2018 2017 Net sales (all credit) Lesss Cost of goods sold Gross profits Less: Other operating expenses Earnings before interest, taxes, depreciation, and S00500 $490 $36 30
media%2F61b%2F61b5173d-4072-47d8-9bed-16
Construct the DuPont ROA and ROE breakdowns for Lake of Egypt Marina, Inc. (Do not round intermediate calculations. Round your answers to 2 decimal places.) DuPont Analysis times a times ! ROA ROE times I =
Answer #1

ROA = Net Income / Total Assets,

It explains the Net income earned by the company by employing the assets of the company.

Under DuPont ROA

ROA = Profit Margin % * Asset Turnover = (Net Profit / Sales) * ( Sales / Total Assets ) = (245 / 700) * (700 / 500)

= 35% * 1.4 times = 49%

ROE = Net Inome / Equity.

It explains the Net income earned for shareholders.

Under DuPont ROE

ROE = (Net Profit / Sales) * ( Sales / Total Assets ) * (Total Assets / Equity )

= (245 / 700) * ( 700 / 500) * ( 500 / 349)

= 35% * 1.4 times * 1.43 times = 70.07%